# Proactive and Reactive Constraint Programming for Stochastic Project Scheduling with Maximal Time-Lags

### v2

@@ 2 The Scheduling Problem
We use the following example from [SFSW13] of a simple RCPSP/max instance. There are five activities: $a$, $b$, $c$, $d$, and $e$, with start times $s_{a}$, $s_{b}$, $s_{c}$, $s_{d}$, and $s_{e}$. The activities have durations of 2, 5, 3, 1, and 2 time units, respectively, and resource requirements of 3, 2, 1, 2, and 2 on a single resource with a capacity of $c_{r}=4$. The precedence relations are as follows: activity $b$ starts at least 2 time units after $a$ starts, activity $b$ starts at least 1 time unit before $c$ starts, activity $c$ cannot start later than 6 time units after $a$ starts, and activity $d$ starts exactly 3 time units before $e$ starts. These constraints can be visualized in a project graph (Figure 1). A feasible solution to this problem is $\{s_{a}=1,s_{b}=3,s_{c}=4,s_{d}=0,s_{e}=3\}$, for which the Gannt chart is visualized in Figure 2.
@@ 3.2.1 Constructing Ordering Constraints
Figure 2: Example Gannt chart (figure adjusted from [FLVX12]). The X-axis is time, and Y-axis shows resource demand. The arrows indicate an example of generating a POS with chaining, starting from a fixed solution schedule.
@@ 6 Conclusion and Future Work
In future work, the same approach could be used to evaluate other scheduling problems, and we can gain more insight into how these methods perform on both well-known problems from the literature and in practical situations. Temporal constraints that are defined from end-to-start can be an interesting problem domain for future work. The $\mathsf{stnu}$ method can extended in multiple directions, e.g. rearranging the order of the jobs online or using probabilistic STNs are both interesting for future work.
@@ Appendix A Reproducibility
We uploaded our code to make it possible to reproduce the results of our research and to facilitate others using a similar benchmarking method. Please refer to the READ.ME of our repository (github.com/kimvandenhouten/AAAI25_SRCPSPmax) to find instructions for requirements and directions to our experiment scripts, and the scripts to run the statistical tests.
@@ B.2 STNU Method
This section gives a more comprehensive explanation of the STNU method. We showcase how the temporal network for the given problem instance looks like and how we can see whether the network is DC.
@@ B.2.2 Step 2: Construct STNU
start afinish astart bfinish bstart cfinish cstart dfinish dstart efinish e2-25-53-3d:1D:-21-1
Figure 7: Step 2a: For each activity two nodes are created, representing its start and end, and Step 2b: We connect the start and end of each activity with a tight link (activities $a$, $b$, $c$, $e$) or a contingent link (activity $d$).
start afinish astart bfinish bstart cfinish cstart dfinish dstart efinish e2-25-53-3d:1D:-21-1-33-2-16
@@
start afinish astart bfinish bstart cfinish cstart dfinish dstart efinish e2-25-53-3d:1D:-21-1-33-2-160000
@@ B.2.3 Step 3: Check DC
The resulting network of Figure 9 does not contain any negative cycles, and therefore the network is dynamically controllable [Mor14]. Note that for this specific example, if we would have used the fixed-point solution and resource chains from Figure 2, we would have found a negative cycle in the resulting STNU and therefore no dynamic controllability, see Figure 10.
start dfinish dstart afinish astart bfinish bstart cfinish cstart efinish e2-25-53-3d:1D:-21-1-33-2-160000-7
Figure 10: Non DC network, due to negative cycle.
@@ B.2.4 Step 4: Run RTE*
The RTE* [Pos22, HP24] algorithm provides the online scheduling strategy that transforms the network into a schedule (i.e. it assigns execution times to each node). The RTE* adapts the execution times to the outcome of the contingent time points, resulting in schedule $\{s_{a}=0,s_{b}=3,s_{c}=4,s_{d}=2,s_{e}=3\}$ if $d_{d}=1$ and $\{s_{a}=0,s_{b}=4,s_{c}=5,s_{d}=2,s_{e}=4\}$ if $d_{d}=2$.
@@ Appendix C Tuning
This section describes the tuning process for the different scheduling methods. First, we investigated the effect of the time limit on solving the deterministic CP. Then, for the proactive approach, we tuned the time limit and the sample selection for the Sample Average Approximation (SAA). For $\mathsf{reactive}$, we tuned the quantile approximation and the time limit for rescheduling.
@@ C.2 Time Limit for CP on Deterministic Instances
For the ubo50 instances PSP1 to PSP10 and the ubo100 instances PSP1 to PSP10, the effect of the time limits of 60 seconds, 600 seconds, and 3600 seconds respectively, is presented in Table 5 and 6. We observe that the solver status does not change when increasing the time limit to 600 seconds, and only flips 1 instance from feasible to infeasible when increasing the time limit from 600 seconds to 3600 seconds. The makespan only improves significantly for the ubo100 instances for higher time limits. In the remaining experiments, we fixed the time limit to 60 seconds for solving deterministic CPs.
@@ C.4 Reactive Approach
Next, we observe the effect of the time limit for rescheduling. We used the same subset of the j10 and ubo50 instances (for both sets $PSP\_1$-$PSP\_20$, used 10 duration samples per instance, and runtime limits of 1, 2, 10 and 30 seconds. The results (in Table 9) are almost similar for the different time limits, this might be because the solver finishes already before the time limit, and the increase in time online has mainly to do with the number of solver calls. In the experimental evaluation, we therefore fixed the rescheduling time limit to $2$ seconds, which we expected to be sufficient for larger or slightly more complicated problems.
@@ D.1 Wilcoxon Test
  • If the two methods have no consistent difference in their relative performances, then the rank-sums should be approximately equal. This is tested with a normal approximation for the Wilcoxon statistic which is outlined by [Cur67]. [Cur67] propose a corrected normal approximation which is needed because of usage of the the Pratt procedure for handling zero differences.
@@ D.2 Z-test for Proportion (Binomial Distribution)
The proportion test investigates whether the is a significant difference between the assumed proportion of wins $p_{0}$ and the observed proportion of wins $p$. In our analysis, two methods are compared and the number of wins for each method is counted based on one metric.
@@ D.3 Magnitude Test
The magnitude test we use is a t-test for two population means, or the method of paired comparisons such as Test 10 in the book by [Kan06]. The test whether there is a significant difference between two population means. The procedure for this paired comparison t-test is as follows:
  • Collect a set of matched pairs (so the results from two different methods on one atomic instance forms a pair).
@@
After the normalization step, it is possible to execute the test with the Scipy package and specifically, its built-in method $scipy.stats.ttest\_rel$.
@@ Appendix E Results
Please find the results of the statistical test in this section. In Section E.1 we discuss the feasibility rations. In Section E.2 we present the visualizations of the partial orderings. In Section E.3, we present the results of the magnitude tests on the double hits that can give some insight into the magnitude differences of the performance metrics between the different methods.
@@ E.1 Feasibility Ratio
First, we analyze the feasibility ratio obtained by the different methods. For $\epsilon=1$, we observe the feasibility ratio obtained by $\mathsf{proactive}_{\mathsf{SAA}}$, $\mathsf{proactive}_{0.9}$, $\mathsf{reactive}$ are the same for instance sets j10 - ubo50, and only for ubo100 $\mathsf{reactive}$ is a little bit lower than the $\mathsf{proactive}_{\mathsf{SAA}}$, $\mathsf{proactive}_{0.9}$. The $\mathsf{stnu}$ method has the lowest feasibility rate but remarkably the difference becomes smaller when the size of the problem grows (the difference is the smallest for ubo100).
@@ E.2.1 Partial Ordering Solution Quality
    • Only for ubo50, all arrows are solid, indicating that for all pairs we found a significant difference according to the Wilcoxon test.
    • For ubo100, there is no significant difference between $\mathsf{proactive}_{\mathsf{SAA}}$ and $\mathsf{proactive}_{0.9}$. Possibly for the larger instances the SAA becomes more difficult to solve, for which reason it is not better than the heuristic $\gamma$-quantile procedure anymore.
@@
    • For j10-j20, the partial ordering is similar to $\epsilon=1$, although for $\epsilon=2$ all arrows are solid (indicating a relation obtained by the Wilcoxon test).
@@ E.2.2 Partial Ordering Time Offline
  • We expected the $\mathsf{proactive}_{0.9}$ and $\mathsf{reactive}$ to be the fastest offline, followed by the $\mathsf{proactive}_{\mathsf{SAA}}$ and $\mathsf{stnu}$, where the partial ordering between the latter two depends on the problem size. When two methods have (almost) only ties (such as $\mathsf{proactive}_{0.9}$ and $\mathsf{reactive}$) not all tests can be executed and a $nan$ will appear.
  • Importantly, the tests are designed in such a way that infeasibilities are also weighted in significance testing. When a method results in an infeasible solution, infinite offline time is assigned to this experiment. For that reason, it can still occur that we find a difference between $\mathsf{reactive}$ and $\mathsf{proactive}_{0.9}$: although they have the same offline procedure, one of the two can still fail which results in infinite offline time.
  • In general, we give preference to the test results of the Wilcoxon test because this test shows a stronger relation than the proportional test, but to fill in the missing information we use the proportion test to test if there is a significant difference in the proportion of wins. We noted, however, that the test outcomes of the two tests can sometimes be contradicting. The Wilcoxon test penalizes infeasibilities more severely, while the proportion test only considers the number of wins. Consequently, it is possible that method 1 produces more infeasible solutions but achieves better metrics, whereas method 2 generates fewer infeasible solutions but worse metrics, resulting in the Wilcoxon test on instances where at least one feasible solution exists tells us method 2 performs better (infeasible solutions are penalized more heavily). Proportion test on instances where at least one feasible solution exists tells us method 1 performs better. The magnitude test on double hits tells us Method 1 performs better as it has better metric values.
@@ E.2.3 Partial Ordering Time Online
  • For $\epsilon=2$ (Figure 16), we find the same pattern as for $\epsilon=1$, for problem sets j10-j30, however, we find that for ubo50 and ubo100, the $\mathsf{stnu}$ starts to outperform other methods, explained by the higher feasibility ratio of the $\mathsf{stnu}$.
@@ E.3 Magnitude Tests
The remarkable things that we observed in the results of the magnitude tests, are highlighted per performance metric in the following subsections.
@@ E.3.1 Magnitude Tests Solution Quality
In Table 14 and Table 15 a comparison of the magnitude differences in solution quality (makespan) on the double hits are provided. Here, the normalized averages indicate how much this difference is. We see that the $\mathsf{stnu}$ method for all pair-wise comparisons on the double hits results in the lowest normalized average of the makespan. In the comparison of $\mathsf{reactive}$ and $\mathsf{proactive}$ we see lower normalized averages for the $\mathsf{reactive}$ approach. Furthermore, we observe that for $\epsilon=2$ the magnitude differences become larger for the different pairs.
@@ E.3.2 Magnitude Tests Time Offline
For time offline, $\epsilon=2$, ubo50 and ubo100, the Wilcoxon test found that the $\mathsf{stnu}$ is better than e.g. $\mathsf{proactive}_{0.9}$ and $\mathsf{reactive}$, while looking at the magnitude results on double hits (this is an important difference in test procedure!), we find that the $\mathsf{proactive}_{0.9}$ and $\mathsf{reactive}$ are much faster offline than $\mathsf{stnu}$, which is also expected. Then, we observe no difference between $\mathsf{proactive}_{0.9}$ and $\mathsf{reactive}$, which can be explained by the fact that both have the same offline procedure and we select only double hits for the test.
@@
j30 $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{react}$-$\mathsf{stnu}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$-$\mathsf{stnu}$ $\mathsf{pro}_{0.9}$-$\mathsf{pro}_{\mathsf{SAA}}$ stnu-$\mathsf{pro}_{\mathsf{SAA}}$
@@
ubo50 $\mathsf{pro}_{0.9}$-$\mathsf{react}$ $\mathsf{react}$-$\mathsf{stnu}$ $\mathsf{react}$-$\mathsf{pro}_{\mathsf{SAA}}$ $\mathsf{pro}_{0.9}$-$\mathsf{stnu}$ $\mathsf{pro}_{0.9}$-$\mathsf{pro}_{\mathsf{SAA}}$ stnu-$\mathsf{pro}_{\mathsf{SAA}}$
@@ E.4 Summary
In our main paper, we employed the statistical analysis from this document to provide a summarizing partial ordering per metric including a brief summary of the main findings. We decide to present the most occuring patterns, and in text highlight any important exceptions.
@@ Appendix F Results Tables Partial Orderings
Table 20: Pairwise comparison on schedule quality (makespan) for noise factor $\epsilon=1$. Using a Wilcoxon test and a proportion test. Including all instances for which at least one of the two methods found a feasible solution. Note that the ordering matters: the first method showed is the better of the two in each paier method 1 - method 2. Each cell shows on the first row [nr pairs] z-value (p-value) of the Wilcoxon test with (*) for p $<$ 0.05. Each cell shows on the second row [nr pairs] z-value (p-value) with (*) for p $<$ 0.05 for the proportion test.
@@
Table 21: Pairwise comparison on schedule quality (makespan) for noise factor $\epsilon=2$. Using a Wilcoxon test and a proportion test. Including all instances for which at least one of the two methods found a feasible solution. Note that the ordering matters: the first method showed is the better of the two in each pair method 1 - method 2. Each cell shows on the first row [nr pairs] z-value (p-value) of the Wilcoxon test with (*) for p $<$ 0.05. Each cell shows on the second row [nr pairs] z-value (p-value) with (*) for p $<$ 0.05 for the proportion test.
@@
Table 22: Pairwise comparison on time offline for noise factor $\epsilon=1$. Using a Wilcoxon test and a proportion test. Including all instances for which at least one of the two methods found a feasible solution. Note that the ordering matters: the first method showed is the better of the two in each pair method 1 - method 2 according to Wilcoxon. Each cell shows on the first row [nr pairs] z-value (p-value) of the Wilcoxon test with (*) for p $<$ 0.05. Each cell shows on the second row [nr pairs] ratio of wins (p-value) with (*) for p $<$ 0.05
@@
Table 23: Pairwise comparison on time offline for noise factor $\epsilon=2$. Using a Wilcoxon test and a proportion test. Including all instances for which at least one of the two methods found a feasible solution. Note that the ordering matters: the first method showed is the better of the two in each pair method 1 - method 2 according to Wilcoxon. Each cell shows on the first row [nr pairs] z-value (p-value) of the Wilcoxon test with (*) for p $<$ 0.05. Each cell shows on the second row [nr pairs] ratio of wins (p-value) with (*) for p $<$ 0.05
@@
Table 24: Pairwise comparison on time online for noise factor $\epsilon=1$. Using a Wilcoxon test and a proportion test. Including all instances for which at least one of the two methods found a feasible solution. Note that the ordering matters: the first method showed is the better of the two in each pair method 1 - method 2. Each cell shows on the first row [nr pairs] z-value (p-value) of the Wilcoxon test with (*) for p $<$ 0.05. Each cell shows on the second row [nr pairs] proportion (p-value) with (*) for p $<$ 0.05
@@
Table 25: Pairwise comparison on time online for noise factor $\epsilon=2$. Using a Wilcoxon test and a proportion test. Including all instances for which at least one of the two methods found a feasible solution. Note that the ordering matters: the first method showed is the better of the two in each paier method 1 - method 2. Each cell shows on the first row [nr pairs] z-value (p-value) of the Wilcoxon test with (*) for p $<$ 0.05. Each cell shows on the second row [nr pairs] proportion (p-value) with (*) for p $<$ 0.05

### v4

@@ 2 The Scheduling Problem
We use the following example from [SFSW13] of a simple RCPSP/max instance. There are five activities: $a$, $b$, $c$, $d$, and $e$, with start times $s_{a}$, $s_{b}$, $s_{c}$, $s_{d}$, and $s_{e}$. The activities have durations of 2, 5, 3, 1, and 2 time units, respectively, and resource requirements of 3, 2, 1, 2, and 2 on a single resource with a capacity of $c_{r}=4$. The precedence relations are as follows: activity $b$ starts at least 2 time units after $a$ starts, activity $b$ starts at least 1 time unit before $c$ starts, activity $c$ cannot start later than 6 time units after $a$ starts, and activity $d$ starts exactly 3 time units before $e$ starts. These constraints can be visualized in a project graph (Figure 1). A feasible solution to this problem is $\{s_{a}=1,s_{b}=3,s_{c}=4,s_{d}=0,s_{e}=3\}$, for which the Gantt chart is visualized in Figure 2.
@@ 3.2.1 Constructing Ordering Constraints
Figure 2: Example Gantt chart (figure adjusted from [FLVX12]). The X-axis is time, and Y-axis shows resource demand. The arrows indicate an example of generating a POS with chaining, starting from a fixed solution schedule.
@@ 6 Conclusion and Future Work
In future work, the same approach could be used to evaluate other scheduling problems, and we can gain more insight into how these methods perform on both well-known problems from the literature and in practical situations. Temporal constraints that are defined from end-to-start can be an interesting problem domain for future work. The $\mathsf{stnu}$ method can extended in multiple directions, e.g., rearranging the order of the jobs online or using probabilistic STNs are both interesting for future work.
@@ Appendix A Reproducibility
We uploaded our code to make it possible to reproduce the results of our research and to facilitate others using a similar benchmarking method. Please refer to the READ.ME of our repository (github.com/kimvandenhouten/AAAI25_SRCPSPmax) to find instructions for requirements and directions for our experiment scripts and the scripts to run the statistical tests.
@@ B.2 STNU Method
This section gives a more comprehensive explanation of the STNU method. We showcase what the temporal network for the given problem instance looks like and how we can see whether the network is DC.
@@ B.2.2 Step 2: Construct STNU
start afinish astart bfinish bstart cfinish cstart dfinish dstart efinish e2-25-53-3d:1D:-22-2
Figure 7: Step 2a: For each activity, two nodes are created, representing its start and end, and Step 2b: We connect the start and end of each activity with a tight link (activities $a$, $b$, $c$, $e$) or a contingent link (activity $d$).
start afinish astart bfinish bstart cfinish cstart dfinish dstart efinish e2-25-53-3d:1D:-22-23-3-2-16
@@
start afinish astart bfinish bstart cfinish cstart dfinish dstart efinish e2-25-53-3d:1D:-22-23-3-2-160000
@@ B.2.3 Step 3: Check DC
The resulting network of Figure 9 does not contain any negative cycles, and therefore the network is dynamically controllable [Mor14]. Note that for this specific example, if we had used the fixed-point solution and resource chains from Figure 2, we would have found a negative cycle in the resulting STNU and, therefore, no dynamic controllability; see Figure 10.
start dfinish dstart afinish astart bfinish bstart cfinish cstart efinish e2-25-53-3d:1D:-22-23-3-2-160000-1
[…]
@@ B.2.4 Step 4: Run RTE*
The RTE* [Pos22, HP24] algorithm provides the online scheduling strategy that transforms the network into a schedule (i.e. it assigns execution times to each node). The RTE* adapts the execution times to the outcome of the contingent time points, resulting in schedule $\{s_{a}=0,s_{b}=2,s_{c}=5,s_{d}=4,s_{e}=7\}$ if $d_{d}=1$ and $\{s_{a}=0,s_{b}=2,s_{c}=6,s_{d}=4,s_{e}=7\}$ if $d_{d}=2$.
@@ Appendix C Tuning
This section describes the tuning process for the different scheduling methods. First, we investigate the effect of the time limit on solving the deterministic CP. Then, for the proactive approach, we tuned the time limit and the sample selection for the Sample Average Approximation (SAA). For $\mathsf{reactive}$, we tuned the quantile approximation and the time limit for rescheduling.
@@ C.2 Time Limit for CP on Deterministic Instances
For the ubo50 instances PSP1 to PSP10 and the ubo100 instances PSP1 to PSP10, the effect of the time limits of 60 seconds, 600 seconds, and 3600 seconds respectively, is presented in Table 5 and 6. We observe that the solver status does not change when increasing the time limit to 600 seconds and only flips 1 instance from feasible to infeasible when increasing the time limit from 600 seconds to 3600 seconds. The makespan only improves significantly for the ubo100 instances for higher time limits. In the remaining experiments, we fixed the time limit to 60 seconds for solving deterministic CPs.
@@ C.4 Reactive Approach
Next, we observe the effect of the time limit for rescheduling. We used the same subset of the j10 and ubo50 instances (for both sets $PSP\_1$-$PSP\_20$, used 10 duration samples per instance, and runtime limits of 1, 2, 10 and 30 seconds. The results (in Table 9) are almost similar for the different time limits, and this might be because the solver finishes already before the time limit, and the increase in time online has mainly to do with the number of solver calls. In the experimental evaluation, we therefore fixed the rescheduling time limit to $2$ seconds, which we expected to be sufficient for larger or slightly more complicated problems.
@@ D.1 Wilcoxon Test
  • If the two methods have no consistent difference in their relative performances, then the rank sums should be approximately equal. This is tested with a normal approximation for the Wilcoxon statistic, which is outlined by [Cur67]. [Cur67] propose a corrected normal approximation, which is needed because of the usage of the Pratt procedure for handling zero differences.
@@ D.2 Z-test for Proportion (Binomial Distribution)
The proportion test investigates whether the is a significant difference between the assumed proportion of wins $p_{0}$ and the observed proportion of wins $p$. In our analysis, two methods are compared, and the number of wins for each method is counted based on one metric.
@@ D.3 Magnitude Test
The magnitude test we use is a t-test for two population means, or the method of paired comparisons such as Test 10 in the book by [Kan06]. The test is whether there is a significant difference between the two population means. The procedure for this paired comparison t-test is as follows:
  • Collect a set of matched pairs (so the results from two different methods on one atomic instance form a pair).
@@
After the normalization step, it is possible to execute the test with the Scipy package and, specifically, its built-in method $scipy.stats.ttest\_rel$.
@@ Appendix E Results
Please find the results of the statistical test in this section. In Section E.1, we discuss the feasibility rations. In Section E.2, we present the visualizations of the partial orderings. In Section E.3, we present the results of the magnitude tests on the double hits that can give some insight into the magnitude differences of the performance metrics between the different methods.
@@ E.1 Feasibility Ratio
First, we analyze the feasibility ratio obtained by the different methods. For $\epsilon=1$, we observe the feasibility ratio obtained by $\mathsf{proactive}_{\mathsf{SAA}}$, $\mathsf{proactive}_{0.9}$, $\mathsf{reactive}$ are the same for instance sets j10 - ubo50, and only for ubo100 $\mathsf{reactive}$ is a little bit lower than the $\mathsf{proactive}_{\mathsf{SAA}}$, $\mathsf{proactive}_{0.9}$. The $\mathsf{stnu}$ method has the lowest feasibility rate, but remarkably, the difference becomes smaller when the size of the problem grows (the difference is the smallest for ubo100).
@@ E.2.1 Partial Ordering Solution Quality
    • Only for ubo50, all arrows are solid, indicating that for all pairs, we found a significant difference according to the Wilcoxon test.
    • For ubo100, there is no significant difference between $\mathsf{proactive}_{\mathsf{SAA}}$ and $\mathsf{proactive}_{0.9}$. Possibly, for the larger instances, the SAA becomes more difficult to solve, for which reason it is not better than the heuristic $\gamma$-quantile procedure anymore.
@@
    • For j10-j20, the partial ordering is similar to $\epsilon=1$, although for $\epsilon=2$, all arrows are solid (indicating a relation obtained by the Wilcoxon test).
@@ E.2.2 Partial Ordering Time Offline
  • We expected the $\mathsf{proactive}_{0.9}$ and $\mathsf{reactive}$ to be the fastest offline, followed by the $\mathsf{proactive}_{\mathsf{SAA}}$ and $\mathsf{stnu}$, where the partial ordering between the latter two depends on the problem size. When two methods have (almost) only ties (such as $\mathsf{proactive}_{0.9}$ and $\mathsf{reactive}$), not all tests can be executed, and a $nan$ will appear.
  • Importantly, the tests are designed in such a way that infeasibilities are also weighted in significance testing. When a method results in an infeasible solution, infinite offline time is assigned to this experiment. For that reason, it can still occur that we find a difference between $\mathsf{reactive}$ and $\mathsf{proactive}_{0.9}$: although they have the same offline procedure, one of the two can still fail, which results in infinite offline time.
  • In general, we give preference to the test results of the Wilcoxon test because this test shows a stronger relation than the proportional test, but to fill in the missing information, we use the proportion test to test if there is a significant difference in the proportion of wins. We noted, however, that the test outcomes of the two tests can sometimes be contradicting. The Wilcoxon test penalizes infeasibilities more severely, while the proportion test only considers the number of wins. Consequently, it is possible that method 1 produces more infeasible solutions but achieves better metrics, whereas method 2 generates fewer infeasible solutions but worse metrics, resulting in the Wilcoxon test on instances where at least one feasible solution exists tells us method 2 performs better (infeasible solutions are penalized more heavily). Proportion test on instances where at least one feasible solution exists tells us method 1 performs better. The magnitude test on double hits tells us Method 1 performs better as it has better metric values.
@@ E.2.3 Partial Ordering Time Online
  • For $\epsilon=2$ (Figure 16), we find the same pattern as for $\epsilon=1$; for problem sets j10-j30. However, we find that for ubo50 and ubo100, the $\mathsf{stnu}$ starts to outperform other methods, explained by the higher feasibility ratio of the $\mathsf{stnu}$.
@@ E.3 Magnitude Tests
The remarkable things that we observed in the results of the magnitude tests are highlighted per performance metric in the following subsections.
@@ E.3.1 Magnitude Tests Solution Quality
In Table 14 and Table 15, a comparison of the magnitude differences in solution quality (makespan) on the double hits is provided. Here, the normalized averages indicate how much this difference is. We see that the $\mathsf{stnu}$ method for all pair-wise comparisons on the double hits results in the lowest normalized average of the makespan. In the comparison of $\mathsf{reactive}$ and $\mathsf{proactive}$ we see lower normalized averages for the $\mathsf{reactive}$ approach. Furthermore, we observe that for $\epsilon=2$ the magnitude differences become larger for the different pairs.
@@ E.3.2 Magnitude Tests Time Offline
For time offline, $\epsilon=2$, ubo50 and ubo100, the Wilcoxon test found that the $\mathsf{stnu}$ is better than e.g. $\mathsf{proactive}_{0.9}$ and $\mathsf{reactive}$ while looking at the magnitude results on double hits (this is an important difference in test procedure!), we find that the $\mathsf{proactive}_{0.9}$ and $\mathsf{reactive}$ are much faster offline than $\mathsf{stnu}$, which is also expected. Then, we observe no difference between $\mathsf{proactive}_{0.9}$ and $\mathsf{reactive}$, which can be explained by the fact that both have the same offline procedure, and we select only double hits for the test.
@@ E.4 Summary
In our main paper, we employed the statistical analysis from this document to provide a summarizing partial ordering per metric including a brief summary of the main findings. We decide to present the most occurring patterns and in the text, highlight any important exceptions.
@@ Appendix F Results Tables Partial Orderings
Table 20: Pairwise comparison on schedule quality (makespan) for noise factor $\epsilon=1$. Using a Wilcoxon test and a proportion test. Including all instances for which at least one of the two methods found a feasible solution. Note that the ordering matters: the first method showed is the better of the two in each pair method 1 - method 2. Each cell shows on the first row [nr pairs] the z-value (p-value) of the Wilcoxon test with (*) for p $<$ 0.05. Each cell shows on the second row [nr pairs] z-value (p-value) with (*) for p $<$ 0.05 for the proportion test.
@@
Table 21: Pairwise comparison on schedule quality (makespan) for noise factor $\epsilon=2$. Using a Wilcoxon test and a proportion test. Including all instances for which at least one of the two methods found a feasible solution. Note that the ordering matters: the first method showed is the better of the two in each pair method 1 - method 2. Each cell shows on the first row [nr pairs] the z-value (p-value) of the Wilcoxon test with (*) for p $<$ 0.05. Each cell shows on the second row [nr pairs] z-value (p-value) with (*) for p $<$ 0.05 for the proportion test.
@@
Table 22: Pairwise comparison on time offline for noise factor $\epsilon=1$. Using a Wilcoxon test and a proportion test. Including all instances for which at least one of the two methods found a feasible solution. Note that the ordering matters: the first method showed is the better of the two in each pair method 1 - method 2 according to Wilcoxon. Each cell shows on the first row [nr pairs] the z-value (p-value) of the Wilcoxon test with (*) for p $<$ 0.05. Each cell shows on the second row [nr pairs] the ratio of wins (p-value) with (*) for p $<$ 0.05.
@@
Table 23: Pairwise comparison on time offline for noise factor $\epsilon=2$. Using a Wilcoxon test and a proportion test. Including all instances for which at least one of the two methods found a feasible solution. Note that the ordering matters: the first method showed is the better of the two in each pair method 1 - method 2 according to Wilcoxon. Each cell shows on the first row [nr pairs] the z-value (p-value) of the Wilcoxon test with (*) for p $<$ 0.05. Each cell shows on the second row [nr pairs] the ratio of wins (p-value) with (*) for p $<$ 0.05.
@@
Table 24: Pairwise comparison on time online for noise factor $\epsilon=1$. Using a Wilcoxon test and a proportion test. Including all instances for which at least one of the two methods found a feasible solution. Note that the ordering matters: the first method showed is the better of the two in each pair method 1 - method 2. Each cell shows on the first row [nr pairs] the z-value (p-value) of the Wilcoxon test with (*) for p $<$ 0.05. Each cell shows on the second row [nr pairs] proportion (p-value) with (*) for p $<$ 0.05.
@@
Table 25: Pairwise comparison on time online for noise factor $\epsilon=2$. Using a Wilcoxon test and a proportion test. Including all instances for which at least one of the two methods found a feasible solution. Note that the ordering matters: the first method showed is the better of the two in each pair method 1 - method 2. Each cell shows on the first row [nr pairs] the z-value (p-value) of the Wilcoxon test with (*) for p $<$ 0.05. Each cell shows on the second row [nr pairs] proportion (p-value) with (*) for p $<$ 0.05.